OpenStudy (anonymous):

if you help you get a medal and a fan solve the system of equations: .5x-.75y=3 -2x+3y=-12

OpenStudy (anonymous):

x will always = 1 so now try it but put x as 1

OpenStudy (anonymous):

its like sqared 1

OpenStudy (uri):

Use the elimination method...mutiply the second eq with 25

OpenStudy (uri):

multiply*

OpenStudy (adilalvi):

multiply eq 2 by 25 it will become -50x+75y=-180..now add eq 1 and 3 as under..5x-75y=3+(-50x+75y=-180) +75y and -75y will cancell each other the remaining x=1.7..now putt 1.7 in eq 1 which will give us y=0.07

OpenStudy (anonymous):

they are equivalent.. 0.o

OpenStudy (uri):

.5x-.75y=3 -2x+3y=-12(multiply by 25) 5x-.75y=3 -50x+75y=-300 ========== -45x=-297 x=-297/-45 x=6.6

OpenStudy (uri):

Now put the value of x in the original eq 1. 5x-.75y=3 5(6.6)-75y=3 Solve it.